# GraphOracle: A Foundation Model for Knowledge Graph Reasoning

### v1

@@ Abstract
Foundation models have demonstrated remarkable capabilities across various domains, but developing analogous models for knowledge graphs presents unique challenges due to their dynamic nature and the need for cross-domain reasoning. To address these issues, we introduce GraphOracle, a relation-centric foundation model that unifies reasoning across knowledge graphs by converting them into Relation-Dependency Graphs (RDG), explicitly encoding compositional patterns with fewer edges than prior methods. A query-dependent attention mechanism is further developed to learn inductive representations for both relations and entities. Pre-training on diverse knowledge graphs, followed by minutes-level fine-tuning, enables effective generalization to unseen entities, relations, and entire graphs. Through comprehensive experiments on 31 diverse benchmarks spanning transductive, inductive, and cross-domain settings, we demonstrate consistent state-of-the-art performance with minimal adaptation, improving the prediction performance by up to 35% compared to the strongest baselines.
@@ 1 Introduction
Foundation Models (FMs) [1], known for their extensive self-supervised pre-training and strong few-shot downstream capabilities, have been extended to graph data via Graph Foundation Models (GFM) [17, 9, 18]. However, GFMs, which primarily focus on general graph structures, encounter limitations when applied to Knowledge Graph (KG) reasoning. This is because KGs uniquely require the explicit modeling of complex associations among their diverse relations—a specialized relational awareness often missing in GFM aggregation frameworks. To bridge this gap, Knowledge Graph Foundation Models (KGFMs) [48, 11, 40] have emerged, specifically adapting the FM concept for multi-relational graphs. KGFMs typically learn relation-invariant representations by pre-training on multiple, diverse KGs. This core process equips them to facilitate robust reasoning over unseen entities, relations, and even entirely new KGs. Consequently, KGFMs endeavor to surmount the persistent challenges intrinsic to conventional KG reasoning methodologies. For instance, Transductive reasoning is limited as it assumes that all entities and relations are known during training [2, 37]. Furthermore, Entity Inductive reasoning [35, 28, 57, 51], while handling new entities by learning rules or paths for reasoning, requires fixed relation sets, thus falling short of fully-inductive reasoning—all limitations KGFMs are designed to address by fostering more adaptable and generalizable reasoning.
KGFMs bring KG reasoning into a Fully-Inductive style, enabling generalization to both new entities and new relations. A key strategy is the construction of abstract graph structures that model interactions between relations, independent of specific vocabularies. For instance, INGRAM [19] constructs a “relation graph” by linking relations that share entities, with edge weights determined by their co-occurrence affinities. ULTRA [11] builds a “meta-graph” from universal relational interaction patterns to further refining the relational interaction paradigm—specifically, four predefined paradigms (head-to-head, head-to-tail, tail-to-head, and tail-to-tail) that model interactions between the conceptual connection points of relations.
(a) Number of edges in different relation graphs.
(b) Grouped performance comparison.
Figure 1: Comparative Analysis of Relation Graph Density and Performance. (a) compares the number of edges in relation graphs generated by different KGFMs, where the scatters denote different benchmark statistics and the curves are fitted with quadratic functions. (b) shows the average MRR performance of different KGFMs across benchmarks grouped by the number of relations.
Despite these advancements and the promise of universal reasoning demonstrated by KGFMs, our analysis of current models reveals significant limitations. Specifically, methods like INGRAM and ULTRA suffer from critical limitations in their relation graph construction process, which generates excessive numbers of edges in the relation graphs, leading to substantial noise and ineffective modeling of relational dependencies. As illustrated in Fig. 1(a), the number of edges in the relation graphs generated by INGRAM and ULTRA grows dramatically as the number of relations in the KG increases. This inefficiency has tangible consequences for model performance. As shown in Fig. 1(b), both INGRAM and ULTRA exhibit performance degradation as the number of relations increases. This suggests that the noise introduced by excessive edges hinders effective generalization across KGs, making it difficult to distinguish genuinely critical compositional patterns among relations. We posit that the bottleneck in fully inductive reasoning is not the sheer number of edges in a relation graph, but the model’s ability to capture the true dependency structure among relations. These findings point to the need for methods that trim superfluous edges while faithfully encoding essential relational dependencies.
@@
Motivated by this insight, we introduce GraphOracle, a relation-centric framework that converts entity–relation interactions into a compact Relation-Dependency Graph (RDG). Unlike INGRAM and ULTRA, which generate excessive connections, our RDG retains only meaningful directed precedence links, yielding significantly fewer edges yet richer compositional patterns, with its effectiveness remaining stable as the number of relations increases (see Fig. 1). To capture relation dependencies for specific query, we design a multi-head attention mechanism that recursively propagates information over the RDG, dynamically assembling relation recipes conditioned on the query. By pre-training on four KGs in the general domain, GraphOracle only needs minimal finetune to achieve exceptional adaptability across transductive, inductive, and cross-domain reasoning tasks, improving performance by over 35% on average compared to state-of-the-art methods. Our key contributions in this work can be summarized as follows:
@@
  • Through comprehensive experiments on 31 diverse benchmarks, we demonstrate that GraphOracle consistently outperforms state-of-the-art methods, particularly in demanding domains such as biomedical KGs where compositional reasoning is critical. We show that the attention weights learned on the RDG can reflect important relation dependencies and reduce noise.
@@ 2 Related Works
2.1 Knowledge Graph Reasoning
Transductive Reasoning. Transductive methods assume that every entity and relation encountered at inference time already appears in the training graph. Embedding models such as ConvE [7], RotatE [33] and DuASE [20] learn fixed low-dimensional representations for all entities and relations, while GNN encoders like R-GCN [30] aggregate neighbourhood information with relation-specific graph convolutions, setting strong baselines under this scenario.
@@ 2.1 Knowledge Graph Reasoning
Entity Inductive Reasoning. Entity inductive KG reasoning relaxes the entity constraint: models must generalise to unseen entities while the relation vocabulary stays fixed. Early solutions leveraged auxiliary cues—text descriptions in content-masking models [31] or ontological features in OntoZSL [13]—and symbolic rule learners such as AMIE [10] and NeuralLP [45]. State-of-the-art approaches employ graph neural networks: GraIL [35] uses enclosing subgraphs, whereas PathCon [39], NBFNet [57], RED-GNN [49], A*Net [56] and AdaProp [50] propagate messages along relational paths to accommodate new entities—yet they still assume known relation types.
Fully-Inductive Reasoning. Fully-inductive settings require models to cope with both unseen entities and unseen relations, posing the most demanding generalisation challenge. RMPI [14] and INGRAM [19] represent relations through auxiliary relation graphs; however, RMPI is limited to subgraph extraction, and INGRAM relies on degree discretisation that hampers transfer to graphs with different relation distributions. Recent works such as ISDEA [12] and MTDEA [54] adopt double-equivariant GNNs to model new nodes and relations simultaneously, but their considerable computational overhead restricts scalability to very large graphs.
2.2 Graph Foundation Models
Foundation Models (FMs) have revolutionized natural language and vision tasks through massive self-supervised pre-training that produces strong zero- and few-shot transfer [1]. Adapting this paradigm to graph-structured data, Graph Foundation Models (GFMs) focus primarily on capturing structural properties and facilitating task-specific graph representations [23]. These GFMs can be grouped into three representative categories: (i) masked-reconstruction GFMs, exemplified by GraphMAE [16], which predicts masked node attributes with a GNN-MLP auto-encoder; (ii) structure-aware Transformer GFMs, such as Graphormer [46], which injects centrality and distance encodings into Transformer layers for whole-graph prediction; and (iii) spectral filter GFMs, like GraphAny [53], which fuses multiple frequency filters via linear attention to perform node classification on arbitrary graphs. Although these GFMs excel on homogeneous or lightly heterogeneous benchmarks, their objectives remain largely structural, leaving them ill-equipped for KG reasoning. Critically, GFMs fail to model the complex associations between different relationship types that exist in KGs, which requires sophisticated relational semantics and logical-rule composition beyond mere structural patterns.
Addressing this gap, Knowledge-Graph Foundation Models (KGFMs) have emerged in several flavours: (i) language-bridged KGFMs (KEPLER [42], KGT5 [29]) that couple textual encoders with triple objectives; (ii) prompt-based inductive KGFMs typified by KG-ICL [5], which unifies inductive GNNs with in-context learning to enable plug-and-play deployment on unseen KGs; and (iii) universal relational KGFMs represented by ULTRA [11] and MEGA [40], which pre-train across multiple KGs to learn relation-invariant embeddings and achieve strong zero-shot reasoning—an avenue our GraphOracle framework seeks to further advance.
@@ 3 Preliminary
A Knowledge Graph (KG) consists of sets of entities $\mathcal{V}$, relations $\mathcal{R}$, and fact triples $\mathcal{F}\subseteq(\mathcal{V}\times\mathcal{R}\times\mathcal{V})$ as $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{F})$. $(e_{q},r_{q},e_{a})$ is a triple in KG where $e_{q},e_{a}\in\mathcal{V}$ and $r_{q}\in\mathcal{R}$.
In transductive setup, the KG at both training and inference time are the same, i.e., $(\mathcal{V}_{tra}=\mathcal{V}_{inf})\land(\mathcal{R}_{tra}=\mathcal{R}_{inf})$, while in entity inductive setup, the entities set $\mathcal{V}$ at training and inference are different, $(\mathcal{V}_{tra}\neq V_{inf})\land(\mathcal{R}_{tra}=\mathcal{R}_{inf})$. In this work, we consider fully-inductive settings where $(\mathcal{V}_{tra}\neq\mathcal{V}_{inf})\land(\mathcal{R}_{tra}\neq\mathcal{R}_{inf})$ for a broader generalization across KGs.
@@
$\bm{h}_{r_{q}}^{\ell}(e_{q},e_{y})=\bigoplus\nolimits_{(e_{x},r,e_{y})\in\mathcal{N}(e_{y})}\!\!\bm{h}_{r_{q}}^{\ell-1}(e_{q},e_{x})\otimes\phi(r,r_{q}),$ (1)
where $\bm{h}_{r_{q}}^{0}(e_{q},e)=\bm{1}$ if $e=e_{q}$, and $\bm{0}$ otherwise, $\oplus$/$\otimes$ / ⊗ are learnable additive and multiplicative operators, and $\phi$ encodes relation-type compatibility. After $L$ steps’ iteration, the answer is ranked by the score $s(e_{q},r_{q},e_{a})=\bm{w}_{s}^{\top}\bm{h}_{r_{q}}^{L}(e_{q},e_{a})$. The learning objective is formulated in a contrastive approach that maximizes the log-likelihood of correct triples in the training set, which amounts to minimizing:
$\mathcal{L}_{\text{train}}=-\!\!\!\!\!\!\!\!\!\!\sum_{(e_{q},r_{q},e_{a})\in\mathcal{F}_{\text{train}}}\bigg{[}\log\sigma\big{(}s(e_{q},r_{q},e_{a})\big{)}+\!\!\!\!\!\!\!\!\!\sum_{e^{\prime}_{n}\in\mathcal{N}_{(e_{q},r_{q},e_{a})}}\!\!\!\log\Big{(}1-\sigma\big{(}s(e_{q},r_{q},e^{\prime}_{n})\big{)}\Big{)}\bigg{]},$ (2)
where $\sigma$ is a sigmoid function. GraphOracle retains this framework and introduces a Relation-Dependency Graph (RDG) pre-training objective that endows $\phi$ and $\bm{w}^{L}$ with universal semantics, enabling zero-shot generalization to both unseen entities and unseen relation vocabularies.
@@ 4 The Proposed Method
Figure 2: Overview of the GraphOracle process for predicting the answer entity $e_{a}$ from a given query $(e_{1},r_{1},?)$: Given a Knowledge Graph, we first construct the Relation Dependency Graph (RDG). Then, a multi-head attention mechanism combined with a GNN is used to propagate messages among RDG to obtain relation representations $\bm{R}_{q}$, which are then used in a second GNN for message passing over entity representations. Finally, a scoring function evaluates candidate entities based on the aggregated entity representations, ranking them for answer entity prediction.
@@
4.1 Relation-Dependency Graph (RDG) Construction
To build an effective KGFM capable of cross KG generalization, we must capture the fundamental dependent patterns through which one relation can be represented by a combination of others. Our key innovation is transforming the entity-relation interaction space into a relation-dependency interaction manifold that explicitly models how relations influence each other.
@@ 4.1 Relation-Dependency Graph (RDG) Construction
$\Phi(\mathcal{F})=\bigcup\nolimits_{e,e^{\prime},e^{\prime\prime}\in\mathcal{V}}\{(r_{i},r_{j})\mid(e,r_{i},e^{\prime})\in\mathcal{F}\land(e^{\prime},r_{j},e^{\prime\prime})\in\mathcal{F}\}.$ (3)
Then, the RDG is defined as $\mathcal{G}^{\mathcal{R}}=(\mathcal{R},\mathcal{E}^{\mathcal{R}})$, where the node set $\mathcal{R}$ contains all the relations and the edge sets $\mathcal{E}^{\mathcal{R}}=\Phi(\mathcal{F})$ includes relation dependencies induced by entity-mediated pathways.
This transformation alters the conceptual framework, shifting from an entity-centric perspective to a relation-interaction manifold where compositional connections between relations become explicit. Each directed edge $(r_{i},r_{j})$ in $\mathcal{G}^{\mathcal{R}}$ represents a potential relation-dependency pathway, indicating that relation $r_{i}$ preconditions relation $r_{j}$ through their sequential interaction mediated by a common entity substrate. The edge structure encodes compositional relational semantics, capturing how one relation may influence the probability or applicability of another relation in sequence.
@@
4.2 Relation Representation Learning on Relation-Dependency Graph

### v2

# GraphOracle: Efficient Fully-Inductive Knowledge Graph Reasoning via Relation-Dependency Graphs

@@ Abstract
Knowledge graph reasoning in the fully-inductive setting—where both entities and relations at test time are unseen during training—remains an open challenge. In this work, we introduce GraphOracle, a novel framework that achieves robust fully-inductive reasoning by transforming each knowledge graph into a Relation-Dependency Graph (RDG). The RDG encodes directed precedence links between relations, capturing essential compositional patterns while drastically reducing graph density. Conditioned on a query relation, a multi-head attention mechanism propagates information over the RDG to produce context-aware relation embeddings. These embeddings then guide a second GNN to perform inductive message passing over the original knowledge graph, enabling prediction on entirely new entities and relations. Comprehensive experiments on 60 benchmarks demonstrate that GraphOracle outperforms prior methods by up to 25% in fully-inductive and 28% in cross-domain scenarios. Our analysis further confirms that the compact RDG structure and attention-based propagation are key to efficient and accurate generalization
@@ 1 Introduction
Knowledge graphs (KGs) encode structured knowledge as entity–relation–entity triples, serving as the backbone for scientific discovery, web-scale reasoning, and intelligent systems (Cai2025FusionKGLLM; Bai2025AutoSchemaKG; Dong2025GLMTripleGen; guo2025decouplingcontinualsemanticsegmentation; lin2025seagentselfevolutiontrajectoryoptimization; du2025graphmaster). The central challenge in KG reasoning is link prediction: given an incomplete graph and a query $(h,r,?)$, predict the missing tail entity $t$ (Du2025Mixture). Inductive KG reasoning requires models to generalize to facts that were not explicitly observed during training. In the most challenging scenarios, models must perform fully-inductive (zhang2025trix; cui2024promptkg) reasoning—handling both unseen entities and relations during inference. Cross-domain generalization (wang2025mdgfm; hong2025stability; lin2025unifiedgnn) further requires reasoning over entirely different knowledge graphs containing 100% novel entities and relations. Both settings pose a fundamental compositional generalization challenge: models must recombine learned relational patterns to infer new facts in completely unfamiliar contexts.
Recent approaches to fully-inductive reasoning, including INGRAM (lee2023ingram) and ULTRA (galkin2024ultra), attempt to address this challenge by constructing auxiliary relation graphs $G_{R}=(R,E_{R})$ that capture transferable relational patterns. However, as knowledge graphs scale, this paradigm exposes two fundamental limitations that severely hinder their effectiveness in real-world applications. (1) These methods rely on co-occurrence statistics to connect relations, creating dense graphs with $|E_{R}|=\Theta(|R|^{2})$ edges that inflate computational costs to $O(|R|^{3}\cdot H)$ for $H$-layer message passing. The proliferation of spurious connections obscures meaningful compositional signals, while symmetric treatment of relation pairs erases the inherent directionality of logical composition. (2) Current models compress each relation into a single fixed embedding, forcing individual representations to capture diverse semantic roles across vastly different query contexts. For instance, the relation ”associated with” may connect proteins to diseases in biomedical contexts but link authors to topics in academic graphs, yet existing methods use the same representation regardless of context.
These observations raise a critical research question: How to design a KG reasoning framework that captures meaningful relational dependencies while maintaining computational efficiency? This requires moving beyond dense co-occurrence graphs to a sparse, directed structure that preserves compositional patterns, and replacing fixed relation embeddings with dynamic, query-dependent representations.
@@
To realize this design, we propose GraphOracle, a relation-centric framework that transforms entity–relation interactions into a compact Relation-Dependency Graph (RDG). Unlike INGRAM and ULTRA, which generate excessive connections, our RDG retains only meaningful directed precedence links, yielding significantly fewer edges yet richer compositional patterns, with its effectiveness remaining stable as the number of relations increases. To capture relation dependencies for specific query, we design a multi-head attention mechanism that recursively propagates information over the RDG, dynamically assembling relation recipes conditioned on the query. By pre-training on four KGs in the general domain, GraphOracle only needs minimal finetune to achieve exceptional adaptability across transductive, inductive, and cross-domain reasoning tasks, improving performance by over 16.8% on average compared to state-of-the-art methods. Our key contributions in this work can be summarized as follows:
@@
  • Extensive experiments across 60 benchmarks show that GraphOracle consistently outperforms SOTA methods, with particularly strong results in both fully-inductive and cross-domain settings, demonstrating its robustness and generalization capability in challenging scenarios.
@@ 2 Related Works
Knowledge Graph Reasoning
A Knowledge Graph (KG) consists of sets of entities $\mathcal{V}$, relations $\mathcal{R}$, and fact triples $\mathcal{F}\subseteq(\mathcal{V}\times\mathcal{R}\times\mathcal{V})$ as $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{F})$. $(e_{q},r_{q},e_{a})$ is a triple in KG where $e_{q},e_{a}\in\mathcal{V}$ and $r_{q}\in\mathcal{R}$. Knowledge graph reasoning encompasses several increasingly challenging settings based on what information is available during training versus inference. In the transductive setting, both entities and relations remain fixed: $(\mathcal{V}_{tra}=\mathcal{V}_{inf})\land(\mathcal{R}_{tra}=\mathcal{R}_{inf})$. This allows models to learn fixed embeddings for all components. The entity-inductive setting introduces unseen entities at inference while keeping relations fixed: $(\mathcal{V}_{tra}\neq\mathcal{V}_{inf})\land(\mathcal{R}_{tra}=\mathcal{R}_{inf})$. Most challenging is the fully-inductive setting where both entities and relations are novel: $(\mathcal{V}_{tra}\neq\mathcal{V}_{inf})\land(\mathcal{R}_{tra}\neq\mathcal{R}_{inf})$. Beyond these, cross-domain reasoning requires transferring to entirely different knowledge graphs with no shared entities or relations, demanding the most robust generalization capabilities.
2.1 Transductive Reasoning
Transductive methods assume all entities and relations at inference have been seen during training, enabling the use of fixed relation embeddings. Models like ConvE (ConvE), RotatE (RotatE) and DuASE (li2024duase) learn low-dimensional relation and entity embeddings directly, while GNN variants such as R-GCN (schlichtkrull2018modeling) implement relation-specific message passing that effectively parameterizing relation influence via learned embedding-like transformations. These embedding-based approaches form strong baselines but fundamentally cannot generalize beyond their training vocabulary.
2.2 Entity Inductive Reasoning
@@
Entity inductive KG reasoning relaxes the entity constraint while maintaining fixed relation embeddings. Early solutions leveraged auxiliary cues—text descriptions in content-masking models (shi2018open) or ontological features in OntoZSL (geng2021ontozsl)—and symbolic rule learners such as AMIE (galarraga2013amie) and NeuralLP (yang2017neural). More recent approaches like DRUM (DRUM) employ differentiable rule chaining, while RLogic (RLogic) uses symbolic rule matching. SOTA GNN-based methods including GraIL (teru2020inductive), PathCon (wang2021pathcon), NBFNet (zhu2021nbfnet), RED-GNN (Zhang2022RelationalDigraph), A*Net (Zhu2023AStarNet) and AdaProp (Zhang2023AdaProp) propagate messages along relational paths to accommodate new entities—yet they still rely on fixed relation embeddings, limiting their applicability to scenarios with novel relations.
2.3 Fully-Inductive Reasoning
Fully-inductive settings demand handling both unseen entities and relations, requiring explicit relation graph structures. RMPI (geng2023relational) and INGRAM (lee2023ingram) pioneer this direction by constructing undirected relation graphs; however, RMPI is limited to subgraph extraction, and INGRAM’s degree discretization hampers transfer across graphs with different relation distributions. ISDEA (gao2023double) and MTDEA (zhou2023ood) adopt double-equivariant GNNs, but their computational overhead restricts scalability. ULTRA (galkin2024ultra) advances this with interaction-conditioned relation graphs that adapt based on query context. TRIX (zhang2025trix) introduces expressive adjacency motifs for richer relation modeling, while KG-ICL (cui2024kgicl) employs prompt-based relation graphs.
Method Ent-Ind. Full-Ind. Cross-Dom. Relation Representation
RotatE & DuASE Relation Embedding
A*Net & AdaProp Relation Embedding
DRUM Differentiable rule chaining
RLogic Symbolic rule matching
INGRAM Undirected RG
ULTRA Interaction-Conditioned RG
TRIX Expressive Adjacency Motifs RG
KG-ICL Prompt RG
GraphOracle Relation-Dependency Graph
Table 1: Comparison of inductive capabilities and relation representations. “RG” is short for “relation graph”.
2.4 Cross-domain Reasoning
Cross-domain KG reasoning represents the frontier of generalization, transferring patterns to graphs with entirely new entities and relations. Early work relied on domain-agnostic logical rules; recent advances leverage pre-trained graph foundation models. MDGFM (wang2025mdgfm) introduces multi-domain contrastive pre-training, while SAMGPT (zhang2025samgpt) demonstrates strong transfer without textual signals. Stability-GNN (hong2025stability) addresses structural shift through adversarial perturbations, and UnifiedGNN (lin2025unifiedgnn) jointly handles multiple inductive settings via relation adapters. RiemannGFM (liu2025riemanngfm) incorporates geometric regularization, while Text-Free MDGPT (li2025mfgpt) and GraphMFM (cheng2024graphmfm) scale cross-domain pre-training through modality-agnostic masked modeling.
As summarized in Table 1, KG reasoning methods progress from fixed relation embeddings (transductive) to rule-based reasoning (entity-inductive) to explicit relation graphs (fully-inductive). While existing fully-inductive methods construct undirected or interaction-conditioned graphs, they remain limited to single-domain scenarios. GraphOracle uniquely introduces directed relation-dependency graphs that capture compositional patterns, enabling the first successful cross-domain generalization.
Figure 1: Overview of the GraphOracle process that predicts the answer entity $e_{a}$ from a given query $(e_{1},r_{1},?)$: Given a Knowledge Graph, we first construct the Relation Dependency Graph (RDG). Then, a multi-head attention mechanism combined with a GNN is used to propagate messages among RDG to obtain relation representations $\bm{h}_{r|r_{q}}^{L_{r}}$, which are then used in another GNN for message passing over entity representations. Finally, the candidate entities are scored and evaluated based on the aggregated entity representations, and then ranked for answer entity prediction.
@@ 3 Preliminary
$\bm{h}_{r_{q}}^{\ell}(e_{q},e_{y})=\bigoplus_{(e_{x},r,e_{y})\in\mathcal{N}(e_{y})}\!\!\bm{h}_{r_{q}}^{\ell-1}(e_{q},e_{x})\otimes\phi(r,r_{q}),$ (1)
where $\bm{h}_{r_{q}}^{0}(e_{q},e)=\bm{1}$ if $e=e_{q}$, and $\bm{0}$ otherwise, $\oplus/\otimes$ are learnable additive and multiplicative operators. $\phi$ encodes relation-type compatibility. After $L$ steps’ iteration, the answer is ranked by the score $s(e_{q},r_{q},e_{a})=\bm{w}_{s}^{\top}\bm{h}_{r_{q}}^{L}(e_{q},e_{a})$. The learning objective is formulated in a contrastive approach that maximizes the log-likelihood of correct triples in the training set, which amounts to minimizing:
$\mathcal{L}_{\text{train}}=-\!\!\!\!\!\!\!\!\!\sum_{(e_{q},r_{q},e_{a})\in\mathcal{F}_{\text{train}}}$ $\Big[\log\sigma\big(s(e_{q},r_{q},e_{a})\big)$ (2)
$\!+$ $\!\!\!\!\!\!\!\!\sum_{e^{\prime}_{n}\in\mathcal{N}_{(e_{q},r_{q},e_{a})}}\!\!\!\!\!\!\!\!\log\big(1-\sigma(s(e_{q},r_{q},e^{\prime}_{n}))\big)\Big],$
where $\sigma$ is a sigmoid function. GraphOracle retains this framework and introduces a Relation-Dependency Graph pre-training objective that endows $\phi$ with universal semantics, enabling zero-shot generalization to both unseen entities and unseen relation vocabularies.
@@ 4 The Proposed Method
4.1 RDG Construction
To build an effective KGFM capable of cross KG generalization, we must capture the fundamental dependent patterns through which one relation can be represented by a combination of others. Our key innovation is reparameterizing entity–relation interactions as a relation-dependency interaction structure that explicitly captures how relations influence each other.
@@ 4.1 RDG Construction
$\Phi(\mathcal{F})=\!\!\!\!\!\!\bigcup\limits_{e,e^{\prime},e^{\prime\prime}\in\mathcal{V}}\!\!\!\!\{(r_{i},r_{j})\mid(e,r_{i},e^{\prime})\in\mathcal{F}\land(e^{\prime},r_{j},e^{\prime\prime})\in\mathcal{F}\}.$
Then, the RDG is defined as $\mathcal{G}^{\mathcal{R}}=(\mathcal{R},\mathcal{E}^{\mathcal{R}})$, where the node set $\mathcal{R}$ contains all the relations and the edge sets $\mathcal{E}^{\mathcal{R}}=\Phi(\mathcal{F})$ includes relation dependencies induced by entity-mediated pathways. This transformation alters the conceptual framework, shifting from an entity-centric perspective to a relation-interaction manifold where compositional connections between relations become explicit. Each directed edge $(r_{i},r_{j})$ in $\mathcal{G}^{\mathcal{R}}$ represents a potential relation-dependency pathway, indicating that relation $r_{i}$ preconditions relation $r_{j}$ through their sequential interaction over a shared entity context. The edge structure encodes compositional relational semantics, capturing how one relation may influence the probability or applicability of another when they occur in sequence.
@@
4.2 Relation Representation Learning on RDG